20 ft Determine the landing location (coordinates), impact speed and time of flight for the launch conditions shown. 4. 80 ft 20 ft

Determine the landing location (in coordinates), the impact speed, and the time of flight for the launch conditions shown.
Answer #1

Angle of projection is heta.

20 ft/s 4 80 ft 2 l sin φ l cos φ- 20 ft-

From the given figure, 5 3 cos θ ,sin θ , cos o =

Initial velocity of projectile u=20ft/s

Horizontal distance traveled by projectile in time t is  x=ucos heta, t ---(1)

Vertical distance traveled by projectile in time t is y = 80 + u sin θt--gt ---(2)

Coordinates of landing location are r = 20 l cosọ, ---(3)

y-lsinó ---(4)

from (1) and (3) ,ucos heta ,t=20+lcosphi

20 * _ t = 20+1(12t - 20)V5 I =---(5)

from (2) and (4)  80+usin heta t-rac{1}{2}gt^2=lsinphi

20*rac{4}{5} t-rac{1}{2}*32.2*t^2=-80+l*rac{1}{sqrt{5}} ( since g=32.2 ft/s )

(16t - 16.1 80)V5 ---(6)

from (4) and (6) (16 t-16.1*t^2+80)sqrt{5}=rac{(12 t-20)sqrt{5}}{2}

-16. It + 10t + 90 = 0Rightarrow t=2.7,s

using t=2,7 s in equation (6) = (16t-16.1 * t2 + 80) VS = 13.04 ft

x=20+lcosphi=20+13.04*rac{2}{sqrt{5}}=31.7,ft

,y=lsinphi=13.04*rac{1}{sqrt{5}}=5.8,ft

Final velocity components are v_x=ucos heta and v_y=usin heta-gt

3 u,- 1 cos θ-20 *--12 ft/s

v_y=usin heta-gt=20*rac{4}{5}-(32.2)(2.7)=-70.94,ft/s

Magnitude of landing velocity (impact speed) v (2+2122707.95 ft/s 2 1/2 _ 71 95

Coordinates of landing location are (z, y) = (31.7.5.8) ft

Time of flight t=2,7 s

impact speed 71.95 ft/s